OpenStudy (anonymous):

Two factory plants are making TV panels. Yesterday, Plant A produced 8000 panels. One percent of the panels from Plant A and 5% of the panels from Plant B were defective. How many panels did Plant B produce, if the overall percentage of defective panels from the two plants was 4% ?

OpenStudy (kropot72):

Let x be the total number of panels produced yesterday by plant B. Total production by plant A + plant B = 8000 + x Total defectives made by plant A + plant B = 0.04(8000 + x) Number of defectives made by plant A = 8000 * 0.01 = 80 Number of defectives made by plant B = 0.05x Now we can equate the number of defectives as follows: 0.04(8000 + x) = 80 + 0.05x ...........................(1) Now you need to solve equation (1) to find the value of x, the total number of panels produced yesterday by plant B.